7. The NPV and payback period Aa Aa What information does the payback period provide? Suppose you are evaluating a project with the expected future cash inflows shown in the following table. Your boss has asked you to calculate the projects net present value (NPV). You dont know the projects initial cost, but you do know the projects regular, or conventional, payback period is 2.50 years. If the projects ~WACC~ is 7%, the projects NPV (rounded to the nearest dollar) is: Year Cash Flo Year 1 $350,000 Year 2 $475,000 Year 3 $450,000 Year 4 $450,000 $322,098 O $362,361 $402,623 $483,148 Which of the following statements indicate a disadvantage of using the regular payback period (not the discounted payback period) for capital budgeting decisions? Check all that apply The payback period does not take the time value of money into account. The payback period does not take the projects entire life into account. The payback period is calculated using net income instead of cash flows.
Answer #1

1. Initial Investment = 350,000 *1 + 475000 *1+ 450000*0.5 = 1,050,000
NPV of Project = 350,000/(1+7%) + 475,000/(1+7%)2 + 450,000/(1+7%)3 + 450,000/(1+7%)4 -1,050,000 = 402,623

2. Option a and option b are correct.
Option c is not correct

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well